Suppose you deposit R10 000 in a bank and the reserve requirement is 0.25. If the banking


system has zero excess reserves, then the total amount of new money (not counting your


deposit) that can be created is:


A. R10 000


B. R30 000


C. R40 000


D. R 2 500


E. R 7 500


1
Expert's answer
2021-12-06T17:44:44-0500

Solution:

The correct answer is C. R40,000

 

Change in money supply = money multiplier "\times" the initial deposit

 

First derive the money multiplier:

MM = 1/Reserve ratio = "\frac{1}{0.25}" = 4

Change in money supply = 4 "\times" 10,000 = 40,000

 

 The total amount of new money = R40,000